# Comparative study of quantum emitter fabrication in wide bandgap materials using localized electron irradiation

### 07 Dec 2023 v2

@@ III.2 Theoretical calculations
For hBN it is known to host a large variety of optically active emitters [45]. The question arises if this is true for mica as well. This would determine if our fabrication method simply does not work on mica or that there are generally not many optically active defects. To calculate the optical signatures, we employed density functional theory (DFT) calculations (See Methods for the computational details). Here, we take into account the spin polarization as, for the allowed optical transition, it should be preserved. Figure 4(a) depicts the electronic structure of the pristine mica with a bandgap of 7.91 eV. Due to this wide gap property, mica should in principle support defect levels similar to other wide bandgap materials just like hBN [45, 46, 47] and diamond [48, 49, 50]. In these materials, single photon emission originates from a transition between two-level states of occupied and unoccupied defect states, unlike TMDs that are some time from the transition between strained valence and conduction bands [51, 52]. As a consequence, we calculate 20 different types of point defects, mostly native and antisite defects but also carbon impurities, since a scanning electron microscope typically bonds carbon to the surface [34]. For this, we aimed to investigate the existence of two-level states localized far away from the valence and conduction bands greater than 1 eV, the so-called deep-lying defect states.
@@
In Figure 4(b), the V defect introduces two-level states between occupied and unoccupied states contributed by oxygen atoms. However, these are shallow levels which have full occupation at room temperature. Considering a carbon dopant at different sites in Figures 4(c) and (d), both cases do not exhibit any two-level states. Although it likely has an unoccupied state at 5 eV, which can accept some excited electrons, this large transition energy is out of reach for our experiments. In addition, we expect that this system would have a low transition rate. Together with the absence of an occupied isolated defect state, both are unlikely to act as a single photon emitter. It is worth noting that all defects seem not to localize at both bottom-most and top-most states. This indicates that the defect in mica has little influence on deep-lying two-level states. All these behaviors have also been observed in the rest of the defects shown in Supplementary S6 except for Al, where its desired two-level states exist. Thus, this implies that mica hosts many optically inactive defects compared to hBN [45, 47]. As a result, it can be concluded that although mica offers a wide bandgap, which in principle can host many defects, electronic states of those doped defects do not appear in the bandgap region, but rather around the valence and conduction bands. In other words, mica is not a feasible material to support single photon emitters, at least not as suitable as hBN.
IV Conclusion
We extensively demonstrate the investigation of quantum emitters through a localized irradiation process and have provided a valuable insights into the effectiveness of this methodology across the various insulating materials. We ventured beyond mica and explored other insulating materials to comprehensively assess the fabrication experimental evidences. Our fabrication methodology yield the fabrication of diffraction limited isolated bright spot with indication of quantum emission but inconclusive. Our localized irradiation worked effective to generate the yellow emitting single photon emitter in hBN samples from different vendors. However, in between these samples we have seen differences in emitter properties, with samples from one vendor being not very photostable. Our extension to 3D insulating systems like SiC and GaN, where no quantum emitters or emitter ensembles were activated, independent of electron beam parameters reveals the material-dependent response to localized irradiation. The theoretical simulation has also been implemented and revealed that the mica is unlikely to support single photon emission due to the lack of two-level defect states.
@@ IV Conclusion
Overall, our findings highlight the intricate relationship between the choice of insulating material and the effectiveness of localized irradiation in creating quantum emitters. While this methodology has proven robust in some materials like hBN, its efficacy varies, presenting both exciting possibilities and challenges in the quest to engineer quantum emitters for diverse applications in quantum technologies and photonics. Further research and optimization efforts are warranted to harness the full potential of this fabrication technique across different material platforms.
@@ Methods
We used a standard Si/SiO substrate (purchased from Microchem) with a 290 nm thermally grown oxide layer. Afterwards, the substrate is processed using electron beam lithography and a metal lift-off process to realize a cross-grid pattern. The cross gird patterned is used to navigate to the target flake during sample fabrication and characterization processes. The bulk crystal for exfoliation is commercially purchased from hq Graphene (Mica muscovite and hBN-1), 2D Semiconductor (hBN-2). Thin mica, hBN-1 and hBN-2 flakes were first mechanically exfoliated onto a polymer sheet (PDMS, Polydimethylsiloxane) using scotch tape. Afterwards a suitable thin flake is selected under bright field optical microscope and then transferred to a cross grid patterned standard Si/SiO substrate. Additionally, we acquired a 1010 mm 4H semi-insulating-type silicon carbide crystal substrate from MSE Supplies. In addition, we obtained a 2-inch undoped n-type gallium nitride template, which was grown on a sapphire (0001) substrate.
@@
The optical investigation of the quantum emitters was conducted using a commercial time-resolved confocal microscope (PicoQuant MicroTime 200). This setup offers four linearly-polarized excitation lasers with wavelengths of 375 nm, 470 nm, 530 nm, and 640 nm, each with pulse lengths ranging from 40 to 90 ps (FWHM), depending on the wavelength. The excitation power used for all measurements was maintained at approximately 30 W, unless specified otherwise. To effectively distinguish the emitted signals from the excitation laser, the detection path of the setup was equipped with a variety of notch filters, including both long-pass and bandpass filters (inserted depending on the specific laser wavelength). Moreover, the setup featured a dual configuration of single photon avalanche diodes (SPADs) positioned in both arms of a 50:50 beamsplitter. This configuration enabled us to measure the second-order correlation function (g) and conduct other relevant photon statistics analyses. For the acquisition of PL maps for each individual flake, the stage was meticulously scanned with a dwell time of 5 ms, and the laser operated at a repetition rate of 20 MHz. The PL signal is collected using a 100 dry immersion objective with high numerical aperture (NA) of 0.9. The data analysis of the correlation function as well as lifetime measurements is performed with the built-in software (that also takes the instrument response function into account). The data acquisition times for these measurements was 1 min per emitter. The spectrometer (Andor Kymera 328i) is attached to one of the exit ports of the optical setup to collect the spectrum of the emitters.
@@
The DFT calculations were carried out using the plane-wave Vienna Ab initio Simulation Package (VASP) [53, 54]. A projector augmented wave (PAW) was employed for treating the core nuclei and valence electrons [55, 56]. A crystal structure of bulk pristine muscovite mica (KAlSiO[OH]) containing 76 atoms from the database [57] was fully optimized with a cutoff energy of 450 eV and the convergence threshold at 10 eV until the convergence force is less than 0.02 eV/Å. A 543 Monkhorst–Pack reciprocal space grid was implemented for Brillouin zone integration [58]. After structural relaxation of the pristine structure, we obtained the lattice parameters at Å, Å, Å, and , corresponding to other calculations and experiments [59, 60, 61]. It is important to note that experimental band gaps are still under debate and vary among 3.60 eV [60], 5.09 eV [39], and 7.85 eV [62]. We then aimed to tackled this issue by implementing the modified screened hybrid density functional of Heyd–Scuseria–Ernzerhof (HSE) with the Hartree-Fock exact exchange () ratio of 0.37 to circumvent the common underestimation from the conventional Perdew–Burke-Ernzerhof (PBE) [63]. For the sake of computation time, we initially relaxed the structure using the common PBE; then, we applied the HSE functional to estimate the bandgap. This in turn yields 7.91 eV, which agrees with the experimental value and another DFT work [64].
@@ Acknowledgements.
This work was funded by the Federal Ministry of Education and Research (BMBF) under grant number 13N16292 and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Projektnummer 445275953. The authors acknowledge support by the German Space Agency DLR with funds provided by the Federal Ministry for Economic Affairs and Climate Action BMWK under grant numbers 50WM2165 (QUICK3) and 50RP2200 (QuVeKS). T.V. and V.D. acknowledge support from the DFG via the SFB 1375 NOA Project C2 (Project number 398816777). This research is part of the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus. The major instrumentation used in this work was funded by the Free State of Thuringia via the projects 2015 FOR 0005 (ACP-FIB) and 2017 IZN 0012 (InQuoSens). C.C. acknowledges a Development and Promotion of Science and Technology Talents Project (DPST) scholarship by the Royal Thai Government. The computational experiments were supported by resources of the Friedrich Schiller University Jena supported in part by DFG grants INST 275/334-1 FUGG and INST 275/363-1 FUGG. S.S. acknowledges funding support by Mahidol University (Fundamental Fund: fiscal year 2023 by National Science Research and Innovation Fund (NSRF)) and from the NSRF via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation (grant number B05F650024). T.S. is supported via the BMBF funding program Photonics Research Germany (‘LPI‐BT1‐FSU’, FKZ 13N15466) and is integrated into the Leibniz Center for Photonics in Infection Research (LPI). The LPI initiated by Leibniz‐IPHT, Leibniz‐HKI, UKJ and FSU Jena is part of the BMBF national roadmap for research infrastructures.References
- Degen et al. [2017] C. Degen, F. Reinhard, and P. Cappellaro, Quantum sensing, Rev. Mod. Phys. 89, 035002 (2017).
- Naomichi [2020] K. H. M. Naomichi, Discrete-time quantum walk on complex networks for community detection, Phys. Rev. Res. 2, 023378 (2020).
- Abasifard et al. [2023] M. Abasifard, C. Cholsuk, R. G. Pousa, A. Kumar, A. Zand, T. Riel, D. K. L. Oi, and T. Vogl, The ideal wavelength for daylight free-space quantum key distribution, arXiv:2303.02106 (2023), arXiv:2303.02106 [quant-ph] .
- Al‐Juboori et al. [2023] A. Al‐Juboori, H. Z. J. Zeng, M. A. P. Nguyen, X. Ai, A. Laucht, A. Solntsev, M. Toth, R. Malaney, and I. Aharonovich, Quantum key distribution using a quantum emitter in hexagonal boron nitride, Adv. Quantum Technol. 6, 10.1002/qute.202300038 (2023).
- Çağlar Samaner et al. [2022] Çağlar Samaner, S. Paçal, G. Mutlu, K. Uyanık, and S. Ateş, Free‐space quantum key distribution with single photons from defects in hexagonal boron nitride, Adv. Quantum Technol. 5, 10.1002/qute.202200059 (2022).
- Aharonovich et al. [2016] I. Aharonovich, D. Englund, and M. Toth, Solid-state single-photon emitters, Nat. Photonics 10, 631 (2016).
- Tran et al. [2016] T. T. Tran, K. Bray, M. J. Ford, M. Toth, and I. Aharonovich, Quantum emission from hexagonal boron nitride monolayers, Nat. Nanotechnol. 11, 37 (2016).
- Senichev et al. [2021] A. Senichev, Z. O. Martin, S. Peana, D. Sychev, X. Xu, A. S. Lagutchev, A. Boltasseva, and V. M. Shalaev, Room-temperature single-photon emitters in silicon nitride, Sci. Adv. 7, 10.1126/sciadv.abj0627 (2021).
- Lohrmann et al. [2015] A. Lohrmann, N. Iwamoto, Z. Bodrog, S. Castelletto, T. Ohshima, T. Karle, A. Gali, S. Prawer, J. McCallum, and B. Johnson, Single-photon emitting diode in silicon carbide, Nat. Commun. 6, 7783 (2015).
- Wang et al. [2018] J. Wang, Y. Zhou, Z. Wang, A. Rasmita, J. Yang, X. Li, H. J. von Bardeleben, and W. Gao, Bright room temperature single photon source at telecom range in cubic silicon carbide, Nat. Commun. 9, 4106 (2018).
- Stachurski et al. [2022] J. Stachurski, S. Tamariz, G. Callsen, R. Butté, and N. Grandjean, Single photon emission and recombination dynamics in self-assembled gan/aln quantum dots, Light Sci. Appl. 11, 114 (2022).
- Exarhos et al. [2019] A. L. Exarhos, D. A. Hopper, R. N. Patel, M. W. Doherty, and L. C. Bassett, Magnetic-field-dependent quantum emission in hexagonal boron nitride at room temperature, Nat. Commun. 10, 222 (2019).
- Huang et al. [2022] M. Huang, J. Zhou, D. Chen, H. Lu, N. J. McLaughlin, S. Li, M. Alghamdi, D. Djugba, J. Shi, H. Wang, and C. R. Du, Wide field imaging of van der waals ferromagnet fe3gete2 by spin defects in hexagonal boron nitride, Nat. Commun. 13, 5369 (2022).
- Vogl et al. [2021] T. Vogl, H. Knopf, M. Weissflog, P. K. Lam, and F. Eilenberger, Sensitive single-photon test of extended quantum theory with two-dimensional hexagonal boron nitride, Phys. Rev. Res. 3, 013296 (2021).
- Chen et al. [2020] Y. Chen, T. N. Tran, N. M. H. Duong, C. Li, M. Toth, C. Bradac, I. Aharonovich, A. Solntsev, and T. T. Tran, Optical thermometry with quantum emitters in hexagonal boron nitride, ACS Appl. Mater. Interfaces. 12, 25464 (2020).
- Gottscholl et al. [2021] A. Gottscholl, M. Diez, V. Soltamov, C. Kasper, D. Krauße, A. Sperlich, M. Kianinia, C. Bradac, I. Aharonovich, and V. Dyakonov, Spin defects in hbn as promising temperature, pressure and magnetic field quantum sensors, Nat. Commun. 12, 4480 (2021).
- Conlon et al. [2023] L. O. Conlon, T. Vogl, C. D. Marciniak, I. Pogorelov, S. K. Yung, F. Eilenberger, D. W. Berry, F. S. Santana, R. Blatt, T. Monz, P. K. Lam, and S. M. Assad, Approaching optimal entangling collective measurements on quantum computing platforms, Nat. Phys. 19, 351 (2023).
- Pfaff et al. [2013] W. Pfaff, T. H. Taminiau, L. Robledo, H. Bernien, M. Markham, D. J. Twitchen, and R. Hanson, Demonstration of entanglement-by-measurement of solid-state qubits, Nat. Phys. 9, 29 (2013).
- Cholsuk et al. [2023a] C. Cholsuk, A. Cakan, S. Suwanna, and T. Vogl, Identifying electronic transitions of defects in hexagonal boron nitride for quantum memories, arXiv:2310.20645v1 (2023a), arXiv:2310.20645v1 [quant-ph] .
- Nateeboon et al. [2023] T. Nateeboon, C. Cholsuk, S. Suwanna, and T. Vogl, Modeling the performance and bandwidth of adiabatic quantum memories, arXiv:2306.07855 (2023), arXiv:2306.07855 [quant-ph] .
- Ruf et al. [2021] M. Ruf, N. H. Wan, H. Choi, D. Englund, and R. Hanson, Quantum networks based on color centers in diamond, J. Appl. Phys. 130, 10.1063/5.0056534 (2021).
- Doherty et al. [2013] M. W. Doherty, N. B. Manson, P. Delaney, F. Jelezko, J. Wrachtrup, and L. C. Hollenberg, The nitrogen-vacancy colour centre in diamond, Phys. Rep. 528, 1 (2013).
- Bradac et al. [2019] C. Bradac, W. Gao, J. Forneris, M. E. Trusheim, and I. Aharonovich, Quantum nanophotonics with group iv defects in diamond, Nat. Commun. 10, 5625 (2019).
- Palacios-Berraquero et al. [2017] C. Palacios-Berraquero, D. M. Kara, A. R.-P. Montblanch, M. Barbone, P. Latawiec, D. Yoon, A. K. Ott, M. Loncar, A. C. Ferrari, and M. Atatüre, Large-scale quantum-emitter arrays in atomically thin semiconductors, Nat. Commun. 8, 15093 (2017).
- Klein et al. [2019] J. Klein, M. Lorke, M. Florian, F. Sigger, L. Sigl, S. Rey, J. Wierzbowski, J. Cerne, K. Müller, E. Mitterreiter, P. Zimmermann, T. Taniguchi, K. Watanabe, U. Wurstbauer, M. Kaniber, M. Knap, R. Schmidt, J. J. Finley, and A. W. Holleitner, Site-selectively generated photon emitters in monolayer mos2 via local helium ion irradiation, Nat. Commun. 10, 2755 (2019).
- Vogl et al. [2019a] T. Vogl, M. W. Doherty, B. C. Buchler, Y. Lu, and P. K. Lam, Atomic localization of quantum emitters in multilayer hexagonal boron nitride, Nanoscale 11, 14362 (2019a).
- Vogl et al. [2017] T. Vogl, Y. Lu, and P. K. Lam, Room temperature single photon source using fiber-integrated hexagonal boron nitride, J. Phys. D 50, 295101 (2017).
- Alexandra [2021] X. M. X. Z. O. S. D. V. L. A. S. C. Y. P. T. T. W. K. S. V. M. B. Alexandra, Creating quantum emitters in hexagonal boron nitride deterministically on chip-compatible substrates., Nano Lett. 21, 8182 (2021).
- Vogl et al. [2018] T. Vogl, G. Campbell, B. C. Buchler, Y. Lu, and P. K. Lam, Fabrication and deterministic transfer of high-quality quantum emitters in hexagonal boron nitride, ACS Photonics 5, 2305 (2018).
- Chejanovsky et al. [2016] N. Chejanovsky, M. Rezai, F. Paolucci, Y. Kim, T. Rendler, W. Rouabeh, F. F. de Oliveira, P. Herlinger, A. Denisenko, S. Yang, I. Gerhardt, A. Finkler, J. H. Smet, and J. Wrachtrup, Structural attributes and photodynamics of visible spectrum quantum emitters in hexagonal boron nitride, Nano Lett. 16, 7037 (2016).
- Vogl et al. [2019b] T. Vogl, K. Sripathy, A. Sharma, P. Reddy, J. Sullivan, J. R. Machacek, L. Zhang, F. Karouta, B. C. Buchler, M. W. Doherty, Y. Lu, and P. K. Lam, Radiation tolerance of two-dimensional material-based devices for space applications, Nat. Commun. 10, 1202 (2019b).
- Fou [2021] Position-controlled quantum emitters with reproducible emission wavelength in hexagonal boron nitride, Nat. Commun. 12, 3779 (2021).
- Gale et al. [2022] A. Gale, C. Li, Y. Chen, K. Watanabe, T. Taniguchi, I. Aharonovich, and M. Toth, Site-specific fabrication of blue quantum emitters in hexagonal boron nitride, ACS Photonics 9, 2170 (2022).
- Kumar et al. [2023a] A. Kumar, C. Cholsuk, A. Zand, M. N. Mishuk, T. Matthes, F. Eilenberger, S. Suwanna, and T. Vogl, Localized creation of yellow single photon emitting carbon complexes in hexagonal boron nitride, APL Mater. 11, 10.1063/5.0147560 (2023a).
- 2dd [2022] Defects in Two-Dimensional Materials (Elsevier, 2022).
- Kumar et al. [2023b] A. Kumar, Çağlar Samaner, C. Cholsuk, T. Matthes, S. Paçal, Y. Oyun, A. Zand, R. J. Chapman, G. Saerens, R. Grange, S. Suwanna, S. Ateş, and T. Vogl, Polarization dynamics of solid-state quantum emitters, (2023b).
- Horder et al. [2022] J. Horder, S. J. White, A. Gale, C. Li, K. Watanabe, T. Taniguchi, M. Kianinia, I. Aharonovich, and M. Toth, Coherence properties of electron-beam-activated emitters in hexagonal boron nitride under resonant excitation, Phys. Rev. Appl. 18, 064021 (2022).
- Frisenda et al. [2020] R. Frisenda, Y. Niu, P. Gant, M. Muñoz, and A. Castellanos-Gomez, Naturally occurring van der waals materials, npj 2D Mater. Appl. 4, 38 (2020).
- Kalita and Wary [2016] J. Kalita and G. Wary, Estimation of band gap of muscovite mineral using thermoluminescence (tl) analysis, Physica B Condens. Matter 485, 53 (2016).
- Castelletto et al. [2013] S. Castelletto, B. C. Johnson, V. Ivády, N. Stavrias, T. Umeda, A. Gali, and T. Ohshima, A silicon carbide room-temperature single-photon source, Nat. Mater. 10.1038/NMAT3806 (2013).
- Meunier et al. [2023] M. Meunier, J. J. Eng, Z. Mu, S. Chenot, V. Brändli, P. D. Mierry, W. Gao, and J. Zúñiga-Pérez, Telecom single-photon emitters in gan operating at room temperature: Embedment into bullseye antennas, Nanophotonics 12, 1405 (2023).
- Neumann et al. [2023] M. Neumann, X. Wei, L. Morales-Inostroza, S. Song, S.-G. Lee, K. Watanabe, T. Taniguchi, S. Götzinger, and Y. H. Lee, Organic molecules as origin of visible-range single photon emission from hexagonal boron nitride and mica, ACS Nano 17, 11679 (2023).
- Yang et al. [2016] J. Yang, Z. Wang, F. Wang, R. Xu, J. Tao, S. Zhang, Q. Qin, B. Luther-Davies, C. Jagadish, Z. Yu, and Y. Lu, Atomically thin optical lenses and gratings, Light Sci. Appl. 5, e16046 (2016).
- Roux et al. [2022] S. Roux, C. Fournier, K. Watanabe, T. Taniguchi, J.-P. Hermier, J. Barjon, and A. Delteil, Cathodoluminescence monitoring of quantum emitter activation in hexagonal boron nitride, Appl. Phys. Lett. 121, 10.1063/5.0126357 (2022).
- Cholsuk et al. [2022] C. Cholsuk, S. Suwanna, and T. Vogl, Tailoring the emission wavelength of color centers in hexagonal boron nitride for quantum applications, Nanomaterials 12, 2427 (2022).
- Cholsuk et al. [2023b] C. Cholsuk, S. Suwanna, and T. Vogl, Comprehensive scheme for identifying defects in solid-state quantum systems, J. Phys. Chem. Lett. 14, 6564 (2023b).
- Sajid et al. [2018] A. Sajid, J. R. Reimers, and M. J. Ford, Defect states in hexagonal boron nitride: Assignments of observed properties and prediction of properties relevant to quantum computation, Phys. Rev. B. 97, 1 (2018).
- Ádám Gali [2019] Ádám Gali, Ab initio theory of the nitrogen-vacancy center in diamond, Nanophotonics 8, 1907 (2019).
- Gupta et al. [2019] S. Gupta, J. H. Yang, and B. I. Yakobson, Two-level quantum systems in two-dimensional materials for single photon emission, Nano Lett. 19, 408 (2019).
- Seo et al. [2017] H. Seo, H. Ma, M. Govoni, and G. Galli, Designing defect-based qubit candidates in wide-gap binary semiconductors for solid-state quantum technologies, Phys. Rev. Mater. 1, 1 (2017).
- Parto et al. [2021] K. Parto, S. I. Azzam, K. Banerjee, and G. Moody, Defect and strain engineering of monolayer wse2 enables site-controlled single-photon emission up to 150 k, Nat. Commun. 12, 10.1038/s41467-021-23709-5 (2021).
- Linhart et al. [2019] L. Linhart, M. Paur, V. Smejkal, J. Burgdörfer, T. Mueller, and F. Libisch, Localized intervalley defect excitons as single-photon emitters in wse2, Phys. Rev. Lett. 123, 10.1103/PhysRevLett.123.146401 (2019).
- Kresse and Furthmüller [1996] G. Kresse and J. Furthmüller, Efficiency of ab-initio total energy calculations for metals and semiconductors using a plane-wave basis set, Comput. Mater. Sci. 6, 15 (1996).
- Kresse and Furthmüller [1996] G. Kresse and J. Furthmüller, Efficient iterative schemes for ab initio total-energy calculations using a plane-wave basis set, Phys. Rev. B 54, 11169 (1996).
- Blöchl [1994] P. E. Blöchl, Projector augmented-wave method, Phys. Rev. B 50, 17953 (1994).
- Kresse and Joubert [1999] G. Kresse and D. Joubert, From ultrasoft pseudopotentials to the projector augmented-wave method, Phys. Rev. B 59, 1758 (1999).
- [57] P. Villars and K. Cenzual, Muscovite (kalsio[oh] mon1) crystal structure: Datasheet from “pauling file multinaries edition – 2022” in springer materials, copyright 2023 Springer-Verlag Berlin Heidelberg & Material Phases Data System (MPDS), Switzerland & National Institute for Materials Science (NIMS), Japan.
- Monkhorst and Pack [1976] H. J. Monkhorst and J. D. Pack, Special points for brillouin-zone integrations, Phys. Rev. B. 13, 5188 (1976).
- Franceschi et al. [2023] G. Franceschi, P. Kocán, A. Conti, S. Brandstetter, J. Balajka, I. Sokolović, M. Valtiner, F. Mittendorfer, M. Schmid, M. Setvín, and U. Diebold, Resolving the intrinsic short-range ordering of k+ ions on cleaved muscovite mica, Nat. Commun. 14, 10.1038/s41467-023-35872-y (2023).
- Mukherjee et al. [2022] J. Mukherjee, M. H. Dalsaniya, S. Bhowmick, D. Bhowmik, P. K. Jha, and P. Karmakar, Band gap engineering of the top layer of mica by organized defect formation, Surf. Interfaces. 33, 10.1016/j.surfin.2022.102283 (2022).
- Wu et al. [2022] S. Wu, W. P. Lee, and P. Wu, Origin of observed narrow bandgap of mica nanosheets, Sci. Rep. 12, 10.1038/s41598-022-06820-5 (2022).
- Davidson and Vickers [1972] A. T. Davidson and A. F. Vickers, The optical properties of mica in the vacuum ultraviolet (1972).
- Freysoldt et al. [2014] C. Freysoldt, B. Grabowski, T. Hickel, J. Neugebauer, G. Kresse, A. Janotti, and C. G. V. D. Walle, First-principles calculations for point defects in solids, Rev. Mod. Phys. 86, 253 (2014).
- Vatti et al. [2016] A. K. Vatti, M. Todorova, and J. Neugebauer, Ab initio determined phase diagram of clean and solvated muscovite mica surfaces, Langmuir 32, 1027 (2016).

### 26 Feb 2024 v4

@@ IV.2 Theoretical calculations
It is known that hBN hosts a large variety of optically active emitters [47]. The question arises if this is true for mica as well. Here, we would determine if our fabrication method does not work on mica or if there are generally not many optically active defects using DFT calculations (See Methods for the computational details). Here, to estimate the optical signatures, we consider the spin polarization as it should be preserved for the allowed optical transition. Figure 4(a) depicts the electronic structure of the pristine mica with a bandgap of 7.91 eV. Due to this wide bandgap property, mica should, in principle, support defect levels similar to other wide bandgap materials like hBN [47, 48, 49] and diamond [50, 51, 52]. In these materials, single photon emission originates from a transition between two-level states of occupied and unoccupied defect states, unlike TMDs that are some time from the transition between strained valence and conduction bands [53, 54]. Consequently, we calculate 20 different point defects, primarily native and antisite defects but also carbon impurities, since an SEM typically bonds carbon to the surface [36]. For this, we aimed to investigate the existence of two-level states localized far away from the valence and conduction bands greater than 1 eV, the so-called deep-lying defect states.
@@
In Figure 4(b), the V defect introduces two-level states between occupied and unoccupied states contributed by oxygen atoms. However, these are shallow levels which have full occupation at room temperature. Considering a carbon dopant at different sites in Figures 4(c) and (d), both cases do not exhibit any two-level states. Although it likely has an unoccupied state at 5 eV, which can accept some excited electrons, this large transition energy is out of reach for our experiments. In addition, we expect that this system will have a low transition rate. Together with the absence of an occupied isolated defect state, both are unlikely to act as a single photon emitter. It is worth noting that all defects seem not to localize at both bottom-most and top-most states. This indicates that the defect in mica has little influence on deep-lying two-level states. All these behaviours have also been observed in the rest of the defects shown in Supplementary S7 except for Al, where its desired two-level states exist. Thus, this implies that mica hosts many optically inactive defects compared to hBN [47, 49]. As a result, it can be concluded that although mica offers a wide bandgap, which in principle can host many defects, the electronic states of those doped defects do not appear in the bandgap region but rather around the valence and conduction bands. In other words, mica is not feasible to support single photon emitters, at least not as suitable as hBN.
V Conclusion
We extensively demonstrate the investigation of quantum emitter fabrication using a localized electron irradiation process and provided valuable insights into the effectiveness of electron irradiation across various other insulating materials. Our findings reveal that localized electron irradiation effectively generates single photon emitters in hBN with an optical transition in the yellow region. However, we observe the photo-stability issue between hBN crystals grown differently, attributing to the different intrinsic defect densities or doping levels. This indicates that emitter generation using electron irradiation relies on activation or modification of the charge state of pre-existing defects.
By testing this process on other materials, we can distinguish between a surface complex that is deposited onto the material during our electron irradiation and such activation of lattice defects. This is in particular important as a recent study using organic solvents has found that (at least some) of the hBN as well as mica emitters to be related to organic molecules adhered to the surface of hBN and mica, respectively [46]. We can rule out this possibility for our hBN defects. In addition, we have found that our method does not activate pre-existing defects in the other investigated materials. This could have several reasons where we can only speculate. Maybe there are simply no intrinsic defects that can be activated with an electron beam. Of course, the electrons do not have enough energy to knock out atoms and produce, e.g., vacancies.
Our experimental evidences on mica are also supported by theoretical DFT simulation. The theoretical findings on mica indicate that it is unlikely to support single photon emitters due to the lack of two-level defect states within its band gap. Therefore, the bright emission observed in the irradiated mica sample may be attributed to the generation of surface complexes, similar to a previous observation [46].
@@ V Conclusion
Overall, our findings highlight the intricate relationship between the choice of insulating material and the effectiveness of localized irradiation in creating quantum emitters. While this methodology has proven robust in some materials like hBN, its efficacy varies. It presents exciting possibilities and challenges in the quest to engineer quantum emitters for diverse applications in quantum technologies and photonics. Further research and optimization efforts are warranted to harness the full potential of this fabrication technique across different material platforms.
@@ Methods
We used a standard Si/SiO substrate (purchased from Microchem) with a 290 nm thermally grown oxide layer. Afterwards, the substrate is processed using electron beam lithography and a metal lift-off process to realize a cross-grid pattern. The cross gird patterned is used to navigate to the target flake during sample fabrication and characterization processes. The bulk crystal for exfoliation is commercially purchased from hq Graphene (Muscovite mica and hBN-1), 2D Semiconductor (hBN-2). Thin mica, hBN-1 and hBN-2 flakes were first mechanically exfoliated onto a polymer sheet (PDMS, Polydimethylsiloxane) using scotch tape. Afterwards a suitable thin flake is selected under bright field optical microscope and then transferred to a cross grid patterned standard Si/SiO substrate. Additionally, we acquired a 1010 mm 4H semi-insulating-type silicon carbide crystal substrate from MSE Supplies. In addition, we obtained a 2-inch undoped n-type gallium nitride template, which was grown on a sapphire (0001) substrate.
@@
The optical investigation of the quantum emitters was conducted using a commercial time-resolved confocal microscope (PicoQuant MicroTime 200). This setup offers four linearly-polarized excitation lasers with wavelengths of 375 nm, 470 nm, 530 nm, and 640 nm, each with pulse lengths ranging from 40 to 90 ps (FWHM), depending on the wavelength. The excitation power used for all measurements was maintained at approximately 30 W, unless specified otherwise. To effectively distinguish the emitted signals from the excitation laser, the detection path of the setup was equipped with a variety of notch filters, including both long-pass and bandpass filters (inserted depending on the specific laser wavelength). Moreover, the setup featured a dual configuration of single photon avalanche diodes (SPADs) positioned in both arms of a 50:50 beam-splitter. This configuration enabled us to measure the second-order correlation function (g) and conduct other relevant photon statistics analyses. For the acquisition of PL maps for each individual flake, the stage was meticulously scanned with a dwell time of 5 ms, and the laser operated at a repetition rate of 20 MHz. The PL signal is collected using a 100 dry immersion objective with a high numerical aperture (NA) of 0.9. The data analysis of the correlation function as well as lifetime measurements is performed with the built-in software (that also takes the instrument response function into account). The data acquisition times for these measurements was 1 min per emitter. The spectrometer (Andor Kymera 328i) is attached to one of the exit ports of the optical setup to collect the spectrum of the emitters.
@@
The DFT calculations were carried out using the plane-wave Vienna Ab initio Simulation Package (VASP) [55, 56]. A projector augmented wave (PAW) was employed for treating the core nuclei and valence electrons [57, 58]. A crystal structure of bulk pristine muscovite mica (KAlSiO[OH]) containing 76 atoms from the database [59] was fully optimized with a cutoff energy of 450 eV and the convergence threshold at 10 eV until the convergence force is less than 0.02 eV/Å. A 543 Monkhorst–Pack reciprocal space grid was implemented for Brillouin zone integration [60]. After structural relaxation of the pristine structure, we obtained the lattice parameters at Å, Å, Å, and , corresponding to other calculations and experiments [61, 62, 63]. It is important to note that experimental band gaps are still under debate and vary among 3.60 eV [62], 5.09 eV [41], and 7.85 eV [64]. We then aimed to tackled this issue by implementing the modified screened hybrid density functional of Heyd–Scuseria–Ernzerhof (HSE) with the Hartree-Fock exact exchange () ratio of 0.37 to circumvent the common underestimation from the conventional Perdew–Burke-Ernzerhof (PBE) [65]. For the sake of computation time, we initially relaxed the structure using the common PBE; then, we applied the HSE functional to estimate the bandgap. This in turn yields 7.91 eV, which agrees with the experimental value [64] and another DFT work [66].
@@ Acknowledgements.
This work was funded by the Federal Ministry of Education and Research (BMBF) under grant number 13N16292 and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Projektnummer 445275953. The authors acknowledge support by the German Space Agency DLR with funds provided by the Federal Ministry for Economic Affairs and Climate Action BMWK under grant numbers 50WM2165 (QUICK3) and 50RP2200 (QuVeKS). T.V. and V.D. acknowledge support from the DFG via the SFB 1375 NOA Project C2 (Project number 398816777). This research is part of the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus. The major instrumentation used in this work was funded by the Free State of Thuringia via the projects 2015 FOR 0005 (ACP-FIB) and 2017 IZN 0012 (InQuoSens). C.C. acknowledges a Development and Promotion of Science and Technology Talents Project (DPST) scholarship by the Royal Thai Government. The computational experiments were supported by resources of the Friedrich Schiller University Jena supported in part by DFG grants INST 275/334-1 FUGG and INST 275/363-1 FUGG. S.S. acknowledges funding support by Mahidol University (Fundamental Fund: fiscal year 2023 by National Science Research and Innovation Fund (NSRF)) and from the NSRF via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation (grant number B05F650024). T.S. is supported via the BMBF funding program Photonics Research Germany (‘LPI‐BT1‐FSU’, FKZ 13N15466) and is integrated into the Leibniz Center for Photonics in Infection Research (LPI). The LPI initiated by Leibniz‐IPHT, Leibniz‐HKI, UKJ and FSU Jena is part of the BMBF national roadmap for research infrastructures.VI Supplementary Information
(1) Thickness measurements of Mica Flake, (2) Electron Irradiation Process, (3) Power-resolved photoluminescence measurements, (4) Spectrum measurements from diffraction-limited spots in mica, (5) Additional data on sample hBN-2, (6) Optical characterization of SiC and GaN, (7) Electronic band structures of various defects in mica.